When I try to solve this, I get the answer: all odd numbers. Please Help!!!
The letters, A, B, C, and D all represent different digits. If AB+CA=DA and AB-CA=A, what digit does D represent?
(A)5 (B)6 (C)7 (D)8 (E)9

Answer by mananth(16949)

AB+CA=DA
B+A=A so B=0
..
AB-CA=A
0-A=A
so A=5
AB = 50
Since 50-CA = A
CA =45
C=4
..
50+45=95
50-45=5
...
Can you now find D? I guess so
m.ananth@hotmail.ca